If a 20-kg object experiences a 30-N force for a duration of 0.20-second, then what is the momentum

change of the object?


1
Expert's answer
2021-11-22T10:12:30-0500

We can find the momentum change of the object as follows:


"\Delta p =F\Delta t=30\ N\times 0.2\ s=6\ \dfrac{kg\times m}{s}."